OpenStudy (glorylove):

I'm having difficulty with changing dimensions of plane and solid figures, and I've been stuck on this question for wayyyy too long. Mario Agnello is opening a new pizza shop in town. He plans to offer a 12-inch diameter pizza and an 18-inch diameter pizza. He has set the price of a 12-inch cheese pizza at $8.00, based on the amount of ingredients needed to cover that size crust. If Mario would like to keep the pricing proportional, what should he charge for an 18-inch cheese pizza? Show and explain your solution.

OpenStudy (anonymous):

$8 dollars for 12 inch $1 for 12/8 inch= 1.5 inch or else you can say... 12 inch for $8 1 inch for $ 8/12 = $ 0.667

OpenStudy (glorylove):

So I'm supposed to add that to the existing $8.00 for the price of the 18-inch pizza?

OpenStudy (anonymous):

yup

OpenStudy (glorylove):

Thank you! Is there any way to check it, though?

OpenStudy (anonymous):

using ratio Area of 12-inch pizza: Area of 18-inch pizza=Price of 12-inch pizza:Price of 18-inch pizza pi 36:pi 81=8:x I removed pi for easier calculation 36/81=8/x

OpenStudy (glorylove):

Alright, well thank you for your help!

OpenStudy (anonymous):

your welcome, i did my best!